# Multihead Finite-State Dimension

### v1

@@ 5 A Hierarchy of Predimensions
For each $i\in\mathbb{Z}^{+}$, let $p_{i}$ denote the $i$th prime number: $p_{1}=2$, $p_{2}=3$, etc. For each $x\in\mathbb{N}$, define the multiplicity of $p_{i}$ in $x$ to be
$\nu_{i}(x)=\max\left\{k\in\mathbb{N}\;\middle|\;p_{i}^{k}\mid x\right\}.$
@@
For each $h\in\mathbb{Z}^{+}$, we construct an $(h+1)$-head finite-state gambler $G$ such that, for all $\epsilon>0$ and all $S\in\mathbf{C}$, the sequence $F_{h+1}(S)$ is in the success set of the $(1-\frac{1}{p_{h+1}}+\epsilon)$-gale of $G$. In each round of $p_{h+1}$ steps, the $i$th head of $G$ advances $p_{i}$ times, such that the trailing heads point to the bits referenced in the next parity calculation. This means that $G$ can predict one out of every $p_{h+1}$ bits with certainty, and it bets uniformly on all other bits. This corresponds to its $(1-\frac{1}{p_{h+1}}+\epsilon)$-gale amassing unbounded capital. A complete proof is in Appendix A. ∎
@@
for all $0\leq i\leq n$. Note that $Y[A]=Y[A\cap\mathbb{N}]$; we will only use integer indices of strings and sequences, but it will sometimes be convenient to consider the real intervals those integers belong to.
Let $b=\lfloor\gamma n\rfloor+1$, where $\gamma\in(0,1)\cap\mathbb{Q}$ is a parameter whose value will be defined in Lemma 5.5. Define $u_{h}=Y[[b,n]]$,
$U=\bigcup_{i=1}^{h-1}[\sigma_{i}b-c,\sigma_{i}n+c]$
@@
We prove the following in Appendix A.
@@
Fix $\gamma$ and $j$ as in Lemma 5.5, let $n\in\mathbb{N}$ be sufficiently large, and let $b=\lfloor\gamma n\rfloor+1$. Then
$K(u_{h}\mid u)\geq(n-b)(1-p_{h+1}^{-d-1})-O(\log n).$
Proof.
$K(u_{h}\mid u)$ $\geq K(u_{h}\mid w)-O(1)$ (8)
$\geq K(v_{j}\mid w)-O(\log n)$ (9)
$\geq(n-b)(1-p_{h+1}^{-d-1})-O(\log n).$ (10)
For (9), we first prove the following claim.
Claim 5.7.
$(u_{h},w)\mapsto v_{j}$ is a computable mapping
Proof.
For each $k\in\mathbb{N}$, let
$V^{(k)}_{j}=\left\{t\in V_{j}\mid\nu_{j}(t)\leq k\right\}$
which can be rearranged to obtain a formula for $Y[r]=Y[s\cdot p_{j}^{e}]$:
In particular, since
$\nu_{j}(t)=\nu_{j}(s)+\ell\leq\nu_{j}(s)+e-1\leq\nu_{j}(r)-1\leq k,$
we have $t\in V_{j}^{(k)}$ in this case. Thus, each bit of $Y$ in (11) can be obtained from $u_{h}$, $w$, or $v_{j}^{(k)}$. By induction, we conclude that the mapping is computable. ∎
Therefore,
$\leq K(u_{h},w)-K(w)+O(\log n)$ (by (5) and Claim 5.7)
$\leq K(u_{h}\mid w)+O(\log n).$ (symmetry of information)
For (10), first, observe that
$=\left|\left\{m\in[b,n]\mid\nu_{h+1}(m)\in[0,d]\right\}\right|$
$\geq(n-b+1)-\left\lceil\frac{n-b+1}{p_{h+1}^{d+1}}\right\rceil$
$\geq(n-b)\left(1-p_{h+1}^{-d-1}\right)-O(1).$
$K(R[A],R[B])\geq K(R[0..m])-O(1)\geq m-O(1).$ (12)
It follows that
$\geq K(R[A]\mid R[B])-O(1)$ ($v_{j}\mapsto R[A]$ is computable)
$\geq K(R[A],R[B])-K(R[B])-O(\log n)$ (symmetry of information)
$=m-K(R[B])-O(\log n)$ (by (12))
$\geq m-|B|-O(\log n)$ ($R[B]$ is computable from $|B|$ bits)
$=|A|-O(\log n)$
$=|V_{j}|-O(\log n).$
@@
Let $h\in\mathbb{Z}^{+}$, $G$ an $h$-FSG, $\varepsilon>0$ a rational constant, and $n\in\mathbb{N}$ sufficiently large. Define $b$, $u$, and $u_{h}$ as above. Let $x=u_{h}[b..n]$ and y=Y[0..b1]y=Y[0..b-1]. If
$K(u_{h}\mid u)\geq(n-b)(1-\frac{\varepsilon}{2})$
then
@@
Proof.
$g(z)=\max_{w\sqsubseteq z}\frac{d_{G}(yw)}{d_{G}(y)},$
$g(z)=\max_{w\sqsubseteq z}\prod_{i=1}^{|w|}2\beta(q_{b+i})(w[i]).$
$K(u_{h}\mid u)$ $\leq K(j,t,q)+O(1)$
$\leq K(j)+K(t)+K(q)+O(1)$
$\leq K(j)+O(1)$
$\leq\log j+O(1),$ (14)
$\sum_{w\in\{0,1\}^{\ell}}\frac{d_{G}(yw)}{d_{G}(y)}=2^{\ell},$
As $j$ cannot exceed this number, we can combine this with (14) to get
$K(u_{h}\mid u)<(1-\varepsilon)(n-b)+O(1),$
@@
Let $Y=F_{h+1}(R)$, where $R$ is a Martin-Löf random sequence and $h\in\mathbb{Z}^{+}$. For all $\epsilon\in\mathbb{Q}\cap(0,1)$ and every $h$-head finite-state gambler $G$, the set
$\left\{d_{G}^{(1-2\epsilon)}(Y[0..n-1])\;\middle|\;n\in\mathbb{N}\right\}$
is bounded by a constant.
Proof.
Now let $n\geq N_{0}$, and let $k$ be such that $N_{k-1}<n\leq N_{k+1}$. Then
$Y[0..n-1]=y_{0}y_{1}\cdots y_{k}Y[N_{k}..n-1].$
$\leq d_{G}(y_{0})2^{\epsilon(|y_{1}|+\cdots|y_{k+1}|)}$ (by Lemma 5.8)
$\leq d_{G}(y_{0})2^{\epsilon N_{k+1}}$
$\leq d_{G}(y_{0})2^{\epsilon n/\gamma}$
$\leq d_{G}(y_{0})2^{2\epsilon n}.$ ($\gamma\in(1/2,1)$)
Hence,
$\leq d_{G}(y_{0}),$
which is a constant. ∎
We now turn back to Theorem 5.1. Letting $\epsilon$ approach 0 and applying Lemma 5.9 shows that
$\dim_{\textup{FS}}^{(h)}(Y)=\operatorname{Dim}_{\textup{FS}}^{(h)}(Y)=1,$
while Lemma 5.4 shows that
@@ 6 Instability of $h$-Head Finite-State Predimensions
We prove this using two variants of our function family $F_{h+1}$:
For each $h\in\mathbb{Z}^{+}$, define the functions $F^{\prime}_{h+1}:\mathbf{C}\to\mathbf{C}$ and $F^{\prime\prime}_{h+1}:\mathbf{C}\to\mathbf{C}$ as follows. For all $S\in\mathbf{C}$, let $F^{\prime}_{h+1}(S)[0]=F^{\prime\prime}_{h+1}(S)[0]=S[0]$ and, for all $q\in\mathbb{N}$ and $0\leq r<p_{h+1}$, let
@@
For any fixed $h$ and $\epsilon>0$, we can construct $h$-FSGs whose $(1-\frac{1}{p_{h+1}}+\epsilon)$-gales succeed, for all $S\in\mathbf{C}$, on either $F^{\prime}_{h+1}(S)$ or $F^{\prime\prime}_{h+1}(S)$, using constructions almost identical to the $(h+1)$-FSG that succeeds on $F_{h+1}(S)$ in Lemma 5.4; essentially, we omit either the first or the last trailing head from that construction. This establishes, for all $S\in\mathbf{C}$, that the singletons $\{F^{\prime}_{h+1}(S)\}$ and $\{F^{\prime\prime}_{h+1}(S)\}$ have both $h$-head finite-state predimension and strong predimension at most $1-\frac{1}{p_{h+1}}$.
@@
At a high level, the argument is that by the obliviousness of the head movements, the heads of any $h$-FSG must move identically on both these sequences, which means that for at least one of these sequences, at least one of the relevant h1h-1 bits in any parity calculation must be missing from the information the gambler can readily access. Much as in Lemmas 5.65.8, and 5.9, we show that this missing information implies that at least one of the sequences can be broken into strings of high conditional Kolmogorov complexity given the trailing heads’ information, which bounds the corresponding gale’s rate of growth on each of these strings and thereby on the entire sequence. Our full proof appears in Appendix B.
7 Stability of Multihead Finite-State Dimension
In this section we prove that, notwithstanding the instability of $h$-head finite-state predimension and strong predimension, multihead finite-state dimension and strong dimension are stable under finite unions. This is because, as we now prove, a multihead finite-state gambler with sufficiently many heads can implement a martingale that approximately averages the martingales of a pair of other multihead finite-state gamblers, thereby succeeding wherever either of the pair would succeed, up to $\epsilon$ error.
@@ 7 Stability of Multihead Finite-State Dimension
Proof.
  • $T=T_{1}\times T_{2}$, $Q=Q_{1}\times Q_{2}\times(\mathbb{Q}_{r}\cap[0,1])$.
  • The initial state is $((t_{1},t_{2}),(q_{1},q_{2},\frac{1}{2}))\in T\times Q$.
  • The initial capital is $c_{0}=\frac{c_{1}+c_{2}}{2}=1$.
$d(wb)=kd(w)\beta(q^{\prime},q^{\prime\prime},\alpha)(b).$
and
where $\tilde{d}_{1}(\lambda)=\tilde{d}_{2}(\lambda)=1$.
We now analyze the relationship between $d$, $d_{1}$, and $d_{2}$. Let
[…]
Now we show that for all $w\in\Sigma^{*}$ and $b\in\Sigma$,
For the first,
$=2kd(w)\alpha\beta_{1}(q^{\prime})(b)$ (def. of $\beta$)
$=k\tilde{d}_{1}(w)\frac{\alpha}{\hat{\alpha}}\beta_{1}(q^{\prime})(b),$
and similarly for the second.
Next, we want to show the following:
$\tilde{d}_{1}(w)$ $\geq(1-2^{1-r})^{|w|}d_{1}(w),$ (15)
$\tilde{d}_{2}(w)$ $\geq(1-2^{1-r})^{|w|}d_{2}(w).$ (16)
$(1-2^{1-r})^{|w|}d_{1}(w)=(1-2^{1-r})^{0}\cdot 1=1.$
$=d_{1}(wb)\times(1-2^{1-r})^{|w|}\frac{\alpha}{\hat{\alpha}}$
$\geq d_{1}(wb)\times(1-2^{1-r})^{|w|}(1-2^{1-r})$
$=d_{1}(wb)\times(1-2^{1-r})^{|w|+1}$
$=d_{1}(wb)\times(1-2^{1-r})^{|wb|}.$
$d(w)$ $=\frac{\tilde{d}_{1}(w)+\tilde{d}_{2}(w)}{2}$
$\geq\frac{(1-2^{1-r})^{|w|}}{2}(d_{1}(w)+d_{2}(w))$
$\geq 2^{-\epsilon|w|}(d_{1}(w)+d_{2}(w)).$
$S^{\infty}\big[d_{G_{1}}^{(s)}\big]\cup S^{\infty}\big[d_{G_{2}}^{(s)}\big]\subseteq S^{\infty}\big[d_{G}^{(s+\epsilon)}\big]\quad\text{and}\quad S_{\textup{str}}^{\infty}\big[d_{G_{1}}^{(s)}\big]\cup S_{\textup{str}}^{\infty}\big[d_{G_{2}}^{(s)}\big]\subseteq S_{\textup{str}}^{\infty}\big[d_{G}^{(s+\epsilon)}\big].\qed$
@@
$\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}(X\cup Y)=\max\{\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}(X),\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}(Y)\}$
@@ 8 Conclusion
To our knowledge, the finite-state dimensions $\dim_{\textup{FS}}$ and $\operatorname{Dim}_{\textup{FS}}$ are the most effective (meaning effectivized in the most restrictive way) of the fractal dimensions that have been investigated. It is not hard to see that the multihead finite-state dimensions $\dim_{\textup{FS}}^{\textup{MH}}$ and $\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}$, while not as effective as the finite-state dimensions, are highly effective in the sense that they are more effective than the polynomial-time dimensions $\dim_{\textup{p}}$ and $\operatorname{Dim}_{\textup{p}}$ of [23, 3].
Other dimensions that are highly effective in the above sense have been investigated. These include pushdown dimensions, Lempel–Ziv dimensions, and perhaps polylogspace dimensions [1, 2, 10, 22, 32, 27]. It will be interesting to see how multihead finite-state dimensions are related to these other highly effective dimensions.
Appendix A Appendix: Lemmas and Proofs from Section 5
@@
First, to guarantee $b>\sigma_{h-1}n+c$, we will require $\gamma>\sigma_{h-1}+\frac{c-1}{n}$. Let $\epsilon=\frac{1-\sigma_{h-1}}{4}$, noting that $\sigma_{h-1}+\epsilon<1$. Choose $n$ so that $\frac{c-1}{n}<\epsilon.$

### v2

@@ 5 A Hierarchy of Predimensions
In this section, we prove our main theorem, a strict hierarchy on predimensions, showing that additional heads enhance the predictive power of multi-head finite-state gamblers.
@@
For each $i\in\mathbb{Z}^{+}$, let $p_{i}$ denote the $i$th prime number: $p_{1}=2$, $p_{2}=3$, etc. For each $x\in\mathbb{N}$, define the multiplicity of $p_{i}$ in $x$ to be $\nu_{i}(x)=\max\big\{k\in\mathbb{N}\;\big|\;p_{i}^{k}\mid x\big\}$.
@@
For each $h\in\mathbb{Z}^{+}$, we construct an $(h+1)$-head finite-state gambler $G$ such that, for all $\epsilon>0$ and all $S\in\mathbf{C}$, the sequence $F_{h+1}(S)$ is in the success set of the $(1-\frac{1}{p_{h+1}}+\epsilon)$-gale of $G$. In each round of $p_{h+1}$ steps, the $i$th head of $G$ advances $p_{i}$ times, such that the trailing heads point to the bits referenced in the next parity calculation. This means that $G$ can predict one out of every $p_{h+1}$ bits with certainty, and it bets uniformly on all other bits. This corresponds to its $(1-\frac{1}{p_{h+1}}+\epsilon)$-gale amassing unbounded capital.
A complete proof is in Appendix A. ∎
@@
for all $i\in\{0,\ldots,n\}$. Note that $Y[A]=Y[A\cap\mathbb{N}]$; we will only use integer indices of strings and sequences, but it will sometimes be convenient to consider the real intervals those integers belong to.
Let $b=\lfloor\gamma n\rfloor+1$, where $\gamma\in(0,1)\cap\mathbb{Q}$ is a parameter whose value will be defined in Lemma 5.5. Define $u_{h}=Y[b..n]$,
$U=\bigcup_{i=1}^{h-1}[\sigma_{i}b-c,\sigma_{i}n+c],$
@@
We prove the following four lemmas in Appendix A. Informally, Lemma 5.5 shows that we can tune parameters such that $u$ excludes one of the $h$ regions relevant to calculating parity bits in the suffix $u_{h}$; Lemma 5.6 shows the algorithmic information consequence that $u_{h}$ has high complexity conditioned on $u$; Lemma 5.7 shows that when this conditional Kolmogorov complexity is high, a gambler cannot win quickly while betting on the suffix $u_{h}$; and Lemma 5.8 applies this reasoning inductively to show that a gambler cannot win much on any prefix of $Y$.
@@
Fix $\gamma$ and $j$ as in Lemma 5.5, let $n\in\mathbb{N}$ be sufficiently large, and let $b=\lfloor\gamma n\rfloor+1$. Then $K(u_{h}\mid u)\geq(n-b)(1-p_{h+1}^{-d-1})-O(\log n)$.
@@
Let $h\in\mathbb{Z}^{+}$, $G$ an $h$-FSG, $\varepsilon>0$ a rational constant, and $n\in\mathbb{N}$ sufficiently large. Define $b$, $u$, and $u_{h}$ as above. Let $x=u_{h}[b..n]$ and y=Y[0..b1]y=Y[0..b-1]. If $K(u_{h}\mid u)\geq(n-b)(1-\frac{\varepsilon}{2})$, then
@@
Let $Y=F_{h+1}(R)$, where $R$ is a Martin-Löf random sequence and $h\in\mathbb{Z}^{+}$. For all $\epsilon\in\mathbb{Q}\cap(0,1)$ and every $h$-head finite-state gambler $G$, the set $\big\{d_{G}^{(1-2\epsilon)}(Y[0..n-1])\;\big|\;n\in\mathbb{N}\big\}$ is bounded by a constant.
With these lemmas, we prove Theorem 5.1. Letting $\epsilon$ approach 0 and applying Lemma 5.8 shows that $\dim_{\textup{FS}}^{(h)}(Y)=\operatorname{Dim}_{\textup{FS}}^{(h)}(Y)=1$, while Lemma 5.4 shows that
@@ 6 Instability of $h$-Head Finite-State Predimensions
We now show that unlike typical fractal dimension notions, $h$-head finite-state predimensions are not stable for fixed $h\geq 2$.
@@
We prove this using two variants of our function family $F_{h+1}$, called $F^{\prime}_{h+1}$ and $F^{\prime\prime}_{h+1}$. Recall that the bit in one out of every $h+1$ positions in a sequence $F_{h+1}(S)$ can be computed as the parity of $h$ previous bits in the sequence. At those same positions in $F^{\prime}_{h+1}(S)$ and $F^{\prime\prime}_{h+1}(S)$, the bit can be computed as the parity of only h1h-1 of the $h$ bits involved in the $F_{h+1}(S)$ calculation. In $F^{\prime}_{h+1}(S)$, it is parity of the first h1h-1 of those $h$ bits, while in $F^{\prime\prime}_{h+1}(S)$, it is the parity of the last h1h-1 of those $h$ bits. Intuitively, these sequences together capture the full complexity of $F_{h+1}(S)$, but each of them individually is well-suited to using only h1h-1 trailing heads.
Formally, for each $h\in\mathbb{Z}^{+}$, define the functions $F^{\prime}_{h+1}:\mathbf{C}\to\mathbf{C}$ and $F^{\prime\prime}_{h+1}:\mathbf{C}\to\mathbf{C}$ as follows. For all $S\in\mathbf{C}$, let $F^{\prime}_{h+1}(S)[0]=F^{\prime\prime}_{h+1}(S)[0]=S[0]$ and, for all $q\in\mathbb{N}$ and $0\leq r<p_{h+1}$, let
@@
For any fixed $h$ and $\epsilon>0$, we can construct a pair of $h$-FSGs whose $(1-\frac{1}{p_{h+1}}+\epsilon)$-gales succeed, for all $S\in\mathbf{C}$, on $F^{\prime}_{h+1}(S)$ and on $F^{\prime\prime}_{h+1}(S)$, respectively, using constructions almost identical to the $(h+1)$-FSG that succeeds on $F_{h+1}(S)$ in Lemma 5.4; essentially, we omit either the first or the last trailing head from that construction. This establishes, for all $S\in\mathbf{C}$, that the singletons $\{F^{\prime}_{h+1}(S)\}$ and $\{F^{\prime\prime}_{h+1}(S)\}$ have both $h$-head finite-state predimension and strong predimension at most $1-\frac{1}{p_{h+1}}$.
@@
At a high level, the argument is that by the obliviousness of the head movements, the heads of any $h$-FSG must move identically on both these sequences, which means that for at least one of these sequences, at least one of the relevant h1h-1 bits in any parity calculation must be missing from the information the gambler can readily access. Much as in Lemmas 5.65.7, and 5.8, we show that this missing information implies that at least one of the sequences can be broken into strings of high conditional Kolmogorov complexity given the trailing heads’ information, which bounds the corresponding gale’s rate of growth on each of these strings and thereby on the entire sequence. Our full proof appears in Appendix B.
7 Stability of Multi-Head Finite-State Dimension
In this section we prove that, notwithstanding the instability of $h$-head finite-state predimension and strong predimension, multi-head finite-state dimension and strong dimension are stable under finite unions. This is because, as we now prove, a multi-head finite-state gambler with sufficiently many heads can implement a martingale that approximately averages the martingales of a pair of other multi-head finite-state gamblers, thereby succeeding wherever either of the pair would succeed, up to $\epsilon$ error. Maintaining this approximate average with finitely many states is delicate; a full proof of the following stability theorem appears in Appendix C.
@@ 7 Stability of Multi-Head Finite-State Dimension
$\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}(X\cup Y)=\max\{\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}(X),\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}(Y)\}.$
@@ 8 Conclusion
To our knowledge, the finite-state dimensions $\dim_{\textup{FS}}$ and $\operatorname{Dim}_{\textup{FS}}$ are the most effective (meaning effectivized in the most restrictive way) of the fractal dimensions that have been investigated. It is not hard to see that the multi-head finite-state dimensions $\dim_{\textup{FS}}^{\textup{MH}}$ and $\operatorname{Dim}_{\textup{FS}}^{\textup{MH}}$, while not as effective as the finite-state dimensions, are highly effective in the sense that they are more effective than the polynomial-time dimensions $\dim_{\textup{p}}$ and $\operatorname{Dim}_{\textup{p}}$ of [27, 3].
Other dimensions that are highly effective in the above sense have been investigated. These include pushdown dimensions, Lempel–Ziv dimensions, and perhaps polylogspace dimensions [1, 2, 10, 26, 38, 32]. It will be interesting to see how multi-head finite-state dimensions are related to these other highly effective dimensions.
It will be noted that our motivating question remains open: A sequence $S\in\Sigma^{\omega}$ is multi-head finite-state random if there is no multi-head finite-state gambler $G$ such that $S\in S^{\infty}[d_{G}]$. Is every sequence with multi-head finite-state dimension 1 multi-head finite-state random?
Appendix A Appendix: Proofs from Section 5
@@
First, to guarantee $b>\sigma_{h-1}n+c$, we will require $\gamma>\sigma_{h-1}+\frac{c-1}{n}$.
Let $\epsilon=\frac{1-\sigma_{h-1}}{4}$, noting that $\sigma_{h-1}+\epsilon<1$. Choose $n$ so that $\frac{c-1}{n}<\epsilon.$
@@
Proof of Lemma 5.6.
$K(u_{h}\mid u)$ $\geq K(u_{h}\mid w)-O(1)$ (5)
$\geq K(v_{j}\mid w)-O(\log n)$ (6)
$\geq(n-b)(1-p_{h+1}^{-d-1})-O(\log n).$ (7)
For (6), we first prove the following claim.
Claim A.1.
$(u_{h},w)\mapsto v_{j}$ is a computable mapping.
Proof. For each $k\in\mathbb{N}$, let
$V^{(k)}_{j}=\left\{t\in V_{j}\mid\nu_{j}(t)\leq k\right\}$
which can be rearranged to obtain a formula for $Y[r]=Y[s\cdot p_{j}^{e}]$:
In particular, since
$\nu_{j}(t)=\nu_{j}(s)+\ell\leq\nu_{j}(s)+e-1\leq\nu_{j}(r)-1\leq k,$
we have $t\in V_{j}^{(k)}$ in this case. Thus, each bit of $Y$ in (8) can be obtained from $u_{h}$, $w$, or $v_{j}^{(k)}$. By induction, we conclude that the mapping is computable. $\triangleleft$
Therefore,
$\leq K(u_{h},w)-K(w)+O(\log n)$ (by (1) and Claim A.1)
$\leq K(u_{h}\mid w)+O(\log n).$ (symmetry of information)
For (7), first, observe that
$=\left|\left\{m\in[b,n]\mid\nu_{h+1}(m)\in[0,d]\right\}\right|$
$\geq(n-b+1)-\left\lceil\frac{n-b+1}{p_{h+1}^{d+1}}\right\rceil$
$\geq(n-b)\left(1-p_{h+1}^{-d-1}\right)-O(1).$
$K(R[A],R[B])\geq K(R[0..m])-O(1)\geq m-O(1).$ (9)
It follows that
$\geq K(R[A]\mid R[B])-O(1)$ ($v_{j}\mapsto R[A]$ is computable)
$\geq K(R[A],R[B])-K(R[B])-O(\log n)$ (symmetry of information)
$=m-K(R[B])-O(\log n)$ (by (9))
$\geq m-|B|-O(\log n)$ ($R[B]$ is computable from $|B|$ bits)
$=|A|-O(\log n)$
$=|V_{j}|-O(\log n).$
Proof of Lemma 5.7.
$g(z)=\max_{w\sqsubseteq z}\frac{d_{G}(yw)}{d_{G}(y)},$
$g(z)=\max_{w\sqsubseteq z}\prod_{i=1}^{|w|}2\beta(q_{b+i})(w[i]).$
$K(u_{h}\mid u)$ $\leq K(j,t,q)+O(1)$
$\leq K(j)+K(t)+K(q)+O(1)$
$\leq K(j)+O(1)$
$\leq\log j+O(1),$ (10)
$\sum_{w\in\{0,1\}^{\ell}}\frac{d_{G}(yw)}{d_{G}(y)}=2^{\ell},$
As $j$ cannot exceed this number, we can combine this with (10) to get
$K(u_{h}\mid u)<(1-\varepsilon)(n-b)+O(1),$
Proof of Lemma 5.8.
Now let $n\geq N_{0}$, and let $k$ be such that $N_{k-1}<n\leq N_{k+1}$. Then
$Y[0..n-1]=y_{0}y_{1}\cdots y_{k}Y[N_{k}..n-1].$
$\leq d_{G}(y_{0})2^{\epsilon(|y_{1}|+\cdots|y_{k+1}|)}$ (by Lemma 5.7)
$\leq d_{G}(y_{0})2^{\epsilon N_{k+1}}$
$\leq d_{G}(y_{0})2^{\epsilon n/\gamma}$
$\leq d_{G}(y_{0})2^{2\epsilon n}.$ ($\gamma\in(1/2,1)$)
Hence,
$\leq d_{G}(y_{0}),$
which is a constant. ∎